OpenStudy (anonymous):

a merchant has coffee worth $20 a pound that she wishes to mix with 70 pounds of coffee worth $90 a pound to get a mixture that is worth $30 a pound. How many pound of coffee should be used.

hero (hero):

Solve using methods for systems of equations

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thanks alot

hero (hero):

Did you solve it yet? What did you get for x and y?

hero (hero):

Total Pounds of Coffee: x + 70 = y Total Cost of Coffee: 20x + 90(70) = 30y x + 70 = y 20x + 6300 = 30y 20(y - 70) + 6300 = 30y 20y - 1400 + 6300 = 30y 6300 - 1400 = (30 - 20)y 4900 = 10y 490 = y